### v1

@@ Abstract
Contrastive learning methods in computer vision typically rely on different views of the same image to form pairs. However, in medical imaging, we often seek to compare entire patients with different phenotypes rather than just multiple augmentations of one scan. We propose harnessing clinically relevant tabular data to identify distinct patient phenotypes and form more meaningful pairs in a contrastive learning framework. Our method uses tabular attributes to guide the training of visual representations, without requiring a joint embedding space. We demonstrate its strength using short-axis cardiac MR images and clinical attributes from the UK Biobank, where tabular data helps to more effectively distinguish between patient subgroups. Evaluation on downstream tasks, including fine-tuning and zero-shot prediction of cardiovascular artery diseases and cardiac phenotypes, shows that incorporating tabular data yields stronger visual representations than conventional methods that rely solely on image augmentations or combined image-tabular embeddings. Furthermore, we demonstrate that image encoders trained with tabular guidance are capable of embedding demographic information in their representations, allowing them to use insights from tabular data for unimodal predictions, making them well-suited to real-world medical settings where extensive clinical annotations may not be routinely available at inference time. The code will be available on GitHub.
Keywords:
Visual Representations Contrastive Learning Cardiac MRI
@@ 1 Introduction
Biobanks provide vast quantity of multimodal medical data that can be leveraged to train medical foundation models. Typically, they include tabular data, such as demographic information and medical history, and imaging data, such as MR or CT scans. However, a combination of imaging and tabular modalities has been underexplored compared to the vast research done on image-text models [18, 3]. Meanwhile, in practice physicians often make clinical decisions based on both images and tabular attributes, as information such as patient demographics can provide crucial insights about the patient health. Specifically in cardiology, information such as patient’s sex, age or smoking status are key factors in assessing the risk of cardiovascular diseases [7, 1], which is the leading cause of death worldwide [17]. Thus, developing methods that can integrate information from both images and tabular data is crucial for improved clinical decision-making. The curated tabular attributes in biobanks enable the development of novel methods that can leverage this information to train visual encoders capable of embedding tabular information within their representations, allowing them to use those insights even in the absence of extensive tabular data at inference.
One of the paradigms that has been successful in unimodal and multimodal training is contrastive learning [10]. In an image-only setting, contrastive learning typically operates by generating two different views of the same image via augmentations [5], defining them as a pair. In the case of image-text methods, the pairs are formed between corresponding samples of the two modalities to train a joint embedding space [14]. These methods are effective in generalizing across a variety of downstream tasks, often outperforming fully supervised approaches, especially when labeled data is scarce. However, representations trained under those settings often exhibit an issue of false negatives [11, 6], as similar samples defined as negative pairs are pushed apart in the representation space. To address this issue in the medical domain, several works have redefined pair sampling strategies. MedCLIP [16] leveraged medical reports to assign soft similarity scores between images and textual descriptions, refining supervision through soft semantic matching. Pai et al. [12] redefined positive and negative pairs in medical imaging by grouping 3D patches containing lesions as positives while treating lesion-free regions as negatives. Hager et al. [8] was the first to integrate medical tabular data and images, combining the settings of SimCLR [5] and CLIP [14] to supervise training of a visual representation using tabular attributes. We extend this work by using the clinical tabular attributes to guide contrastive pair selection for unimodal training, eliminating the need for a joint embedding. Recent studies suggest that unimodal training of medical visual representations can outperform language-based supervision [13], motivating our unimodal approach.
Figure 1: A schematic view of TGV. In a batch, the image pairs are defined according to their tabular similarity. The loss function is then calculated between the current sample and all its pairs to train a clinically meaningful representation space.
The contributions of this paper are twofold. First, we introduce Tabular Guide Vision (TGV), a novel contrastive learning training paradigm harnessing tabular data to construct clinically meaningful pairs. Rather than enforcing a shared representation between modalities or relying on image augmentations, we use tabular similarity to construct multiple positive and negative pair assignments within images in a unimodal contrastive learning setting. Second, we propose an approach to overcome the limitation of image-only representations, specifically their lack of zero-shot capabilities, by enabling zero-shot prediction within a visual-only setting. Instead of relying on predefined textual prompts, as in an image-text setting, we use a representative sample of images from the training set as a reference. The prediction is then derived as the mean of the tabular attribute values associated with the most similar image embedding in the reference set. We demonstrate the effectiveness of our method by training a visual cardiac representation using MR images and tabular data from the UK Biobank [15] and evaluate it by predicting cardiac health outcomes and phenotype. Our results demonstrate the potential of TGV to leverage multimodal datasets for unimodal prediction.
@@ 2 Methodology
The novelty of the method is to leverage a multimodal dataset of image-tabular attribute pairs to learn a representation where clinically similar patients are brought closer in the embedding space, guided by medical tabular data (Fig. 1). We follow the basic setting of SimCLR [5], with a vision encoder $E$ encoding the images into embedding $v\in\mathbb{R}^{d}$, followed by a projection head $f_{v}$ mapping the embeddings to $z\in\mathbb{R}^{p}$, where $d$ and $p$ are the dimensions of the embeddings.
2.1 Defining Similarity Pairs
In a batch of size $N$, each image $x_{i}$ is associated with a number of continuous and categorical attributes included in the tabular data $a_{i}=\{a_{con_{i}},a_{cat_{i}}\}$, which are used to calculate a similarity matrix $S\in\mathbb{R}^{N\times N}$.
Given a matrix of normalized continuous variables $a_{con}\in\mathbb{R}^{N\times M}$, where $M$ is the number of the attributes, we calculate a pairwise Euclidean distance matrix $D\in\mathbb{R}^{N\times N}$. The distance is then converted to a similarity and normalized to match the range of categorical similarity, resulting in a continuous data similarity matrix $S_{con}$.
Categorical attributes from the batch are represented in a matrix $a_{cat}\in\mathbb{R}^{N\times B}$, where $B$ is the number of categorical attributes in the dataset. Categorical similarity $S_{cat}\in R^{N\times N}$ is computed via cosine similarity using $a_{cat}$. The continuous and categorical similarities are then used to calculate the combined similarity between the samples in the batch using:
@@ 2.1 Defining Similarity Pairs
where $\lambda$ is a hyperparameter weighting the tabular guidance during the training process towards continuous or categorical attributes.
@@ 2.2 Tabular Data-Guided Visual Learning
Additionally, a threshold $h$ is introduced such that any other image in the batch with a similarity score within $\max(S_{ij})-h$ is also considered a valid pair for $x_{i}$. This approach ensures that all highly similar instances contribute to the contrastive learning process, leveraging intra-batch relationships. With all the pairs in the batch constructed, the contrastive loss function aligns the representations such that the paired ones are pulled closer to each other, while the unpaired ones are pulled away from one another. The loss is defined as:
$L=\frac{1}{N}\sum_{i=1}^{N}-\log\left(\frac{\sum_{j\in\text{positive}}\exp\left(\frac{\langle z_{i},z_{j}\rangle}{\tau}\right)}{\sum_{j=1}^{N}\exp\left(\frac{\langle z_{i},z_{j}\rangle}{\tau}\right)}\right),$ (3)
[…]
2.3 Zero-Shot Prediction
The setting of our proposed unimodal zero-shot prediction is similar to those presented in image-text joint representation spaces [14]. However, instead of computing similarities between images and text embeddings, we make predictions based on similarities between images in the learned embedding space.
@@ 3 Experiments
3.0.1 Dataset
We use the UK Biobank [15] population study (App. Num. XXX) for the training and evaluation of our method. The dataset contains 49,737 pairs of short-axis cardiac MR and tabular data, out of which 39,975 are used for training, 2,794 for validation, and 6,968 for testing. We use volumes of the cardiac MR as the input to our image encoder consisting of 11 slices over 10 frames, uniformly sampled from the initial 50 frames. The images are zero-padded to a square and then cropped to a size 128x128. Our tabular data consists of 24 attributes, including 10 categorical and 14 continuous variables. The full list of the attributes will be available in our GitHub repository. The continuous attributes consist of the cardiac phenotype, e.g., left and right ventricular ejection fraction. The categorical data consist of demographics information, including sex and smoking, and attributes on the presence of CAD disorders split into those diagnosed pre- and post-MR. The attributes include four different CAD (angina pectoris, acute myocardial infarction, other acute ischaemic heart disease, and chronic ischaemic heart disease) following ICD-10 definition as described by Hager et al. [8]. Due to the low frequency of subjects with CAD, we use a disease-balanced sample for the training of a multi-label CAD prediction comprising 6,426 samples. We only consider diseases reported before the scan date for the classification task. For the fine-tuning of the cardiac phenotype predictions, we use a sample of 5,000 images, with the label quality check performed as described by Bai et al. [2]. Further, a disease-balanced sample of 2,000 patients is included in our reference set $P$, used for the zero-shot prediction tasks. The zero-shot predictions are based on at most 20% of the most similar embeddings in the reference set for CAD classification and 2.5% for cardiac phenotype prediction. The final percentage is defined by empirical tuning on the validation set.
3.0.2 Implementation
Table 1: Downstream task performance comparison for multi-label CAD classification (CAD, $\uparrow$) and cardiac phenotype prediction (remaining columns, $\downarrow$). ZS stands for zero-shot, FT for fine-tuning, and Aug for image augmentations.
@@ 3.0.2 Implementation
Training Model CAD LVEF LVEDM LVEDV RVEF RVEDV MYOESV
- Mean-Guess - 4.81 17.85 29.54 4.73 26.73 17.85
Supervised ResNet50 [9] 65.61 4.31 5.59 10.27 3.81 8.55 6.44
SimCLR [5] 61.98 4.72 12.92 23.1 4.72 21.67 13.05
ZS MMCL [8] 61.79 4.4 8.61 15.04 4.48 13.54 9.37
TGV (w/Aug) 66.6 4.1 10.06 16.26 4.18 15.99 10.43
TGV (Ours) 67.97 4.08 7.65 13.7 3.98 12.38 8.16
SimCLR [5] 71.68 3.49 5.25 9.82 3.18 8.25 6.06
FT MMCL [8] 72.91 3.27 5.63 9.95 3.12 7.75 5.51
TGV (w/Aug) 75.31 3.11 4.88 9.29 3.03 7.47 5.28
TGV (Ours) 76.1 3.18 4.86 9.23 2.95 7.39 5.2

### v2

@@ Abstract
Contrastive learning methods in computer vision typically rely on augmented views of the same image or multimodal pretraining strategies that align paired modalities. However, these approaches often overlook semantic relationships between distinct instances, leading to false negatives when semantically similar samples are treated as negatives. This limitation is especially critical in medical imaging domains such as cardiology, where demographic and clinical attributes play a critical role in assessing disease risk and patient outcomes. We introduce a tabular-guided contrastive learning framework that leverages clinically relevant tabular data to identify patient-level similarities and construct more meaningful pairs, enabling semantically aligned representation learning without requiring joint embeddings across modalities. Additionally, we adapt the k-NN algorithm for zero-shot prediction to overcome the lack of zero-shot capability in unimodal representations. We demonstrate the strength of our methods using a large cohort of short-axis cardiac MR images and clinical attributes, where tabular data helps to more effectively distinguish between patient subgroups. Evaluation on downstream tasks, including fine-tuning, linear probing, and zero-shot prediction of cardiovascular artery diseases and cardiac phenotypes, shows that incorporating tabular data guidance yields stronger visual representations than conventional methods that rely solely on image augmentation or combined image-tabular embeddings. Further, we show that our method can generalize to natural images by evaluating it on a car advertisement dataset. The code will be available on GitHub upon acceptance.
@@ 1 Introduction
Figure 1: Contrastive learning methods typically assume a one-to-one correspondence between positive pairs during training, an assumption present in both image augmentation and tabular supervision (similar to CLIP [31], but with tabular data instead of text) approaches. However, that can result in treating clinically similar patients as negative pairs, leading to false negatives.
@@
Biobanks provide a vast quantity of multimodal medical data that can be leveraged to train medical foundation models. Typically, they include tabular data, such as demographic information and medical history, and imaging data, such as magnetic resonance (MR) or computed tomography (CT) scans. However, methods using images and tabular data still remain underexplored [36]. Meanwhile, in practice physicians often make clinical decisions based on both images and tabular attributes, as information such as patient demographics can provide crucial insights about the patient’s health. Specifically in cardiology, attributes such as patient’s sex, age or smoking status are key factors in assessing the risk of cardiovascular diseases [12, 1], which are the leading cause of death worldwide [39]. Thus, developing methods that can integrate information from both images and tabular data is critical for improved clinical decision-making.
Given this clinical challenge, contrastive learning has emerged as a powerful tool for multimodal data integration. In particular, multimodal contrastive learning using language supervision has been vastly explored as a way to train strong vision encoders [43, 3]. This approach was successfully extended to image–tabular data in [18], enabling the pretraining of image encoders supervised by tabular attributes and effectively leveraging the rich clinical information available in biobanks. However, the approach incorporated the standard rigid one-to-one correspondence between training samples used in contrastive learning, overlooking the semantic similarity between instances. As illustrated in Fig. 1, relying solely on image augmentations or image-tabular data pairs in contrastive learning can result in false negatives [24], where clinically similar patients with the same diseases or phenotypes are incorrectly treated as dissimilar. This highlights the need for methods that can leverage tabular information to define clinically informed pair sampling strategies for contrastive learning.
While biobanks provide structured tabular data ideal for training, such information is often unavailable in routine clinical practice, where time constraints and workflow pressures limit comprehensive data collection [11]. Therefore, there is a need for methods that use tabular data during training to produce image encoders capable of accurate unimodal prediction at inference. Although multimodal training has gained popularity as a way to produce strong unimodal image encoders, it remains unclear whether such complexity is necessary. Recent studies show that unimodal training of visual representations can outperform language-based supervision in both natural images [14] and medical settings [30]. Motivated by those findings, we adapt a vision-centric approach, innovatively using tabular data as a proxy to guide the training of visual representations.
To this end, our contributions are as follows:
  • We introduce Tables Guide Vision (TGV), a novel contrastive learning paradigm harnessing tabular data to construct clinically meaningful pair sampling strategy. Rather than enforcing a shared representation between modalities or relying on image augmentation, we use tabular similarity to construct multiple clinically-informed positive and negative pair assignments within images in a unimodal contrastive learning setting.
  • We propose an approach to enable zero-shot prediction from unimodal visual representations by adapting the k-nearest neighbors (k-NN) algorithm. While k-NN aggregation is well-established in machine learning, it has not been systematically applied as an approach for unimodal zero-shot prediction in visual representations. We modify k-NN for this setting by using a representative set of labeled examples for each target class, against which query representations are compared. This representative set enables reliable mean label aggregation across neighbors, allowing for robust zero-shot prediction.
  • We demonstrate the potential of TGV to leverage multimodal datasets for unimodal prediction by training a visual cardiac representation using MR images and tabular data from the UK Biobank [33] and evaluate it by predicting cardiac health outcomes and phenotype.
  • We show that TGV generalizes beyond cardiology by conducting experiments on the DVM [23] dataset, highlighting the potential of our method in other domains.
2 Related Works
Figure 2: Comparison of the proposed tabular guidance (TGV: Tables Guide Vision) approach against other contractive learning approaches. Instead of relying on different views of the same image for training, as in image augmentation approaches, TGV defines pairs between different subjects in the batch based on their tabular similarity. Tabular guidance operates without the need for a joint embedding space between images and tabular data, which is the case in tabular supervision.
2.1 Contrastive Learning
Contrastive learning is a self-supervised representation learning approach that enables models to differentiate between similar and dissimilar samples by comparing pairs of instances. The objective is to align the representations of positive pairs, such as augmented views of the same image, while pushing apart the embeddings of negative pairs, which are typically other instances in the batch [17]. By optimizing a contrastive loss function, such as InfoNCE [28], the model learns an embedding space where semantically similar samples are clustered together, while dissimilar samples are separated. This paradigm has demonstrated strong performance in computer vision, giving rise to a range of influential methods, including SimCLR [6], BYOL [16], and Barlow Twins [40], among others [22].
2.2 Multimodal Contrastive Learning
The success of contrastive methods in vision, motivated their extension to multimodal data, facilitating alignment between text and images and enabling strong zero-shot performance. The first formulation of this paradigm was introduced in the medical domain with ConVIRT [41], which employs a dual-stream architecture composed of a separate image and text encoder. Using a bidirectional contrastive objective, ConVIRT learns a joint representations from paired medical image-text datasets. Subsequently, this approach was adapted and scaled to natural images in CLIP [31], which utilizes large-scale web-scraped image-text pairs and demonstrated zero-shot results competitive with fully supervised models on standard benchmarks. Beyond image-text alignment, contrastive learning has also been extended to other combinations of modalities, such as image-genetics [35, 42] or audio-video.[25, 27].
2.3 Multimodal Image-Tabular Learning
Multimodal learning approaches has also been studied for image-tabular data. MMCL [18] was the first to explore the concept of tabular data supervision in contrastive learning, leveraging rich tabular data to train a strong image encoder. The method uses images and tabular data for training, but is capable of unimodal predictions, similar to CLIP [31]. In contrast, subsequent work has largely focused on using both data modalities at inference. TIP [9] introduced a self-supervised framework for image-tabular pretraining that is robust to missing or incomplete tabular data at inference. However, the method was not evaluated in a fully unimodal setting where tabular data is entirely absent during inference. STiL [10] proposed a semi-supervised, task-specific training framework that combines image and tabular modalities, but it relies on specialized training for each application and does not employ a general-purpose pretrained encoder capable of generalizing across multiple tasks.
2.4 Pairs in Contrastive Learning
A key limitation of contrastive learning is that it overlooks semantic relationships between samples by assuming a rigid one-to-one correspondence between paired instances. As a result, semantically similar samples get pushed away from the anchor, leading to a problem of false negatives [24]. The issue is commonly addressed by redefining the pair assignments to account for inter-instance similarities. SoftCLIP [15] adjusts the training objective by incorporating soft targets defined by fine-grained intra-modal similarity. FFF [4] assigns multiple positive pairs by considering image-text, image-image, and text-text similarities. On the unimodal side, [13] proposed using language similarity between text descriptions to sample semantically similar image pairs for contrastive visual representation learning.
The issue of false negatives has also been studied in the medical domain, where patients may exhibit clinical similarities, such as shared diseases. To address this, several works redefine their pair sampling strategies to incorporate for clinical similarities at training. MedCLIP [37] uses clinical concepts information, while NB-CLIP [38] harnesses medical knowledge graphs to assign soft similarity scores between images and text. In a unimodal setting, [29] defines contrastive pairs in medical imaging by grouping 3D patches containing lesions as positives, while treating lesion-free regions as negatives. In [21], longitudinal patient metadata is used to guide pair selection for retinal OCT, though the approach relies on dataset-specific metadata and predefined rules, limiting its generalizability.
These prior works highlight the limitations posed by rigid pair definitions and the importance of incorporating semantic similarities into contrastive learning. While existing work has explored this challenge in vision-language approaches, the role of tabular data in guiding pair selection has remained largely unexamined. To the best of our knowledge, TGV is the first approach that can use large-scale tabular information to guide contrastive pair assignment between images of different subjects, providing a semantically-informed pairing strategy without introducing additional supervision or relying on handcrafted rules.
@@ 3 Methodology
Tables Guide Vision leverages a multimodal dataset of image-tabular attributes pairs to learn a representation where clinically similar patients are aligned in the embedding space, guided by medical tabular data. Fig. 2 shows how our method differs from image augmentation and tabular supervision based approaches. TGV does not rely on image augmentations, but rather defines the alternative views of the image as different, but clinically similar, subjects in the batch. In comparison with tabular supervision, our approach does not use a joint embedding space. The tabular data is used as a measure of similarity between the patients and is not used as input to any model.
3.1 Vision Encoder
TGV is based on one vision encoder, following the basic setting of SimCLR [6]. The vision encoder $E$ encodes the images into embedding $v\in\mathbb{R}^{d}$, followed by a projection head $f_{v}$ mapping the embeddings to $z\in\mathbb{R}^{p}$, where $d$ and $p$ are the dimensions of the embeddings.
$v=E(x)$ (1)
$z=f_{v}(v),$ (2)
where $x$ represents the input image. The projection head is only used for the training.
3.2 Tabular Similarity
In a batch of size $N$, each image $x_{i}$ is associated with a number of continuous and categorical attributes included in the tabular data $a_{i}=\{a_{con_{i}},a_{cat_{i}}\}$, which are used to calculate a similarity matrix $S\in\mathbb{R}^{N\times N}$. As continuous and categorical features posses different mathematical properties their similarities are calculated separately and then combined into a single similarity matrix.
Categorical attributes from the batch are represented in a matrix $a_{cat}\in\mathbb{R}^{N\times B}$, where $B$ is the number of categorical attributes in the dataset after applying bipolar (-1/1) encoding. Categorical similarity $S_{cat}\in R^{N\times N}$ is computed via cosine similarity using $a_{cat}$.
Given a matrix of normalized continuous variables $a_{con}\in\mathbb{R}^{N\times M}$, where $M$ is the number of the attributes, we calculate a pairwise Euclidean distance matrix $D\in\mathbb{R}^{N\times N}$. The distance is then converted to a similarity and normalized to match the range of categorical similarity, resulting in a continuous data similarity matrix $S_{con}$. The continuous and categorical similarities are then used to calculate the combined similarity between the samples in the batch using:
@@ 3.2 Tabular Similarity
where $\lambda$ is a hyperparameter weighting the tabular guidance towards continuous or categorical attributes.
@@ 3.3 Tabular Data-Guided Visual Learning
Additionally, a threshold $h$ is introduced to allow for sampling of multiple positive pairs, as many subjects in the batch can resemble high tabular similarity. For each anchor image $x_{i}$, let the maximum similarity score to any other image in the batch be:
$M_{i}=\max_{j\neq i}S_{ij}$ (5)
Then, the set of valid positive pairs for $x_{i}$ can be defined as,
$\mathcal{P}_{i}=\left\{x_{j}\,\middle|\,j\neq i,\;S_{ij}\geq M_{i}-h\right\}$ (6)
This approach ensures that all highly similar instances contribute to the contrastive learning process, leveraging intra-batch relationships. With all the pairs in the batch constructed, the contrastive loss function aligns the representations such that the paired ones are pulled closer to each other, while the unpaired ones are pulled away from one another. The loss is defined as:
$L=\frac{1}{N}\sum_{i=1}^{N}-\log\left(\frac{\sum_{j\in\mathcal{P}_{i}}\exp{(\langle z_{i},z_{j}\rangle/\tau)}}{\sum_{j=1}^{N}\exp{(\langle z_{i},z_{j}\rangle/\tau)}}\right),$ (7)
where $\langle\cdot,\cdot\rangle$ stands for cosine similarity, $\tau$ is a temperature parameter scaling the logits.
3.4 Unimodal Zero-Shot Prediction
The setting of our proposed unimodal zero-shot prediction is based on k-NN. It can be considered similar to the zero-shot approach presented in image-text joint representation spaces [31]. However, instead of computing similarities between images and text embeddings, we make predictions based on similarities between images in the learned embedding space.
@@ 4 Experiments
CAD $\uparrow$ LVEF $\downarrow$ LVEDM $\downarrow$ LVEDV $\downarrow$ RVEF $\downarrow$ RVEDV $\downarrow$ MYOESV $\downarrow$
Model ZS FT ZS FT ZS FT ZS FT ZS FT ZS FT ZS FT
Mean-Guess - - 4.81 - 17.85 - 29.54 - 4.73 - 26.73 - 17.85 -
Supervised
ResNet50 [19] - 65.61 - 4.31 - 5.59 - 10.27 - 3.81 - 8.55 - 6.44
Image Augmentation
SimCLR [6] 62.05 71.68 4.72 3.49 13.01 5.25 23.26 9.82 4.71 3.18 21.89 8.25 13.48 6.06
BYOL [16] 56.99 67.32 4.92 3.99 16.43 5.74 27.81 10.31 4.95 3.39 26.29 7.99 16.19 5.96
SimSiam [8] 57.01 69.89 4.93 3.93 16.04 6.59 27.46 11.09 4.94 3.43 25.62 8.52 16.14 5.89
Barlow Twins [40] 55.12 65.01 4.90 3.57 16.54 6.12 27.17 11.01 4.97 3.39 25.98 8.35 16.52 6.14
Tabular Supervision
MMCL [18] 62,49 72.91 4.48 3.27 8.74 5.63 15.12 9.95 4.55 3.12 13.70 7.47 9.54 5.51
Tabular Guidance
TGV (Ours) 68.70 76.1 4.08 3.18 7.64 4.86 13.63 9.23 3.98 2.95 12.43 7.39 8.18 5.2
Table 1: Downstream task performance comparison for multi-label CAD classification measured using AUC and cardiac phenotype prediction (remaining columns, $\downarrow$) using MAE. ZS stands for zero-shot, FT for fine-tuning. The best result is shown in bold, while the second-best is underlined.
4.1 Dataset
We use the UK Biobank [33] population study for the training and evaluation of our method. The dataset contains 49,737 pairs of short-axis cardiac MR and tabular data, out of which 39,975 are used for training, 2,794 for validation, and 6,968 for testing. We use volumes of the cardiac MR as the input to our image encoder consisting of 11 slices over 10 frames, uniformly sampled from the initial 50 frames. The images are zero-padded to a square and then cropped to a size $\times128$. Our tabular data consists of 24 attributes, including 10 categorical and 14 continuous variables (see supplementary materials (SM) Table 1 for full list). The continuous attributes consist of the cardiac phenotype, e.g., left and right ventricular ejection fraction. The categorical data consist of demographics information, including sex and smoking, and attributes on the presence of CAD disorders split into those diagnosed pre- and post-MR. The attributes include four different CAD (angina pectoris, acute myocardial infarction, other acute ischemic heart disease, and chronic ischemic heart disease) following ICD-10 definition as described in [18]. Due to the low frequency of CAD (8% of the dataset), we construct a disease-balanced training sample by including all CAD-positive patients and an equal number of healthy controls, yielding 6,426 training samples for multilabel CAD prediction. We only consider diseases reported before the scan date for the classification task. For the fine-tuning of the cardiac phenotype predictions, we use a sample of 5,000 images, with the label quality check performed as described in [2]. Further, three non-overlapping, disease-balanced samples of 2,000 patients are considered as the representative set $P$, used for the zero-shot prediction tasks. The zero-shot predictions are based on at most 20% of the most similar embeddings in the reference set for CAD classification and 2.5% for cardiac phenotype prediction. The final percentage is defined by empirical tuning on the validation set (see SM Sec. 3.3).
4.2 Implementation